### v2

# arXiv:2602.24153v2[math.AG] 02 Mar 2026

Abstract.
Let $f$ be a defining polynomial of a reduced projective curve $C\subset{\bf P}^{2}$ having only weighted homogeneous singularities. We show that the strong monodromy conjecture for $f$ follows rather easily from arxiv:1609.04801v11 using a formula of Denef and Loeser for the Newton-nondegenerate case with three variables (which can be deduced in the applied case from the one for the two variable case) together with known results about the strong monodromy conjecture in the two variable case. Here an amazing cancellation occurs so that possible counterexamples fail.
Introduction
Let $f$ be a defining polynomial of a reduced projective curve $C\,{\subset}\,{\mathbb{P}}^{2}$ having only weighted homogeneous singularities. Set $d\,{=}\,\deg f$. We have the following
Theorem 1 ([Sa 25, Theorem 4]).
If there is no nonzero vector field of degree 0 annihilating $f$, then k/d-k/d is a root of the Bernstein-Sato polynomial $b_{f}(s)$ of $f$ for any integers $k$ in $[3,d]$.
We can verify moreover the following.
Theorem 2 ([Sa 25, Theorem 5]).
If there is a nonzero vector field of degree 0 annihilating $f$ with d> 4d\,{>}\,4, then $f$ is extremely degenerated, that is, there is a nonzero linear function $\ell$ on ${\mathbb{R}}^{3}$ with integral coefficients such that $f$ is a $\mathbb{C}$}-linear combination of monomials $x^{i}y^{j}z^{k}$ with $\ell(i,j,k)\,{=}\,0$ after a linear coordinate change of ${\mathbb{C}}^{3}$.
In the case d= 3,4d\,{=}\,3,4, the existence of such a vector field implies that either the conclusion of Theorem 2 is satisfied or the curve $C$ is a union of a line and a conic or two conics both intersecting at one point so that 3/d-3/d is a root of the Bernstein-Sato polynomial at a singular point of $C$, see [Sa 25, Remark 8.5]. It is then easy to reduce the strong monodromy conjecture  for the local topological zeta function  $Z^{\rm top}_{f,0}(s)$ (claiming that a pole of $Z^{\rm top}_{f,0}(s)$ is a root of the Bernstein-Sato polynomial $b_{f}(s)$, see [DeLo 92], [BaVe 26]) to the following (by using known results in the two variable case [Lo 88], see Remark 3.1 below).
Theorem 3.
If the conclusion of Theorem 2 holds, then any pole of the local topological zeta function $Z^{\rm top}_{f,0}(s)$ is a pole of the local topological zeta function of a singular point of $C$.
This can be verified rather easily if one admits a formula for the local topological zeta function of a Newton-nondegenerate polynomial of three variables (which is quite amazing), see [DeLo 92]. In the applied case it is possible to deduce this formula from the one for the two variable case (which seems much easier to prove) using the blowup at the origin followed by the ones with line centers together with some computations of rational functions by a computer algebra program such as Macaulay2 or Singular among others. Here it is very surprising that the pole 3/d-3/d disappears  when one calculates $Z^{\rm top}_{f,0}(s)$ even if the Euler number of ${\mathbb{P}}^{2}\,{\setminus}\,C$ does not vanish so that one could expect a counterexample, see Remark 3.2 below.
Note finally that the strong monodromy conjecture has been shown quite recently in the not necessarily reduced curve case under the hypothesis that the associated reduced curve has only weighted homogeneous singularities, see [BaVe 26], where the vanishing of the residue at 3/d-3/d is proved in the needed case. Unfortunately it does not seem to be mentioned there that the following assertion actually holds.
Proposition 1.
Let $Z\,{\subset}\,{\mathbb{P}}^{n-1}$ be a reduced projective hypersurface defined by a polynomial $f$ with $n\,{\geqslant}\,4$. If the hypersurface $Z$ has only homogeneous isolated singularities which are not ordinary double points, then the strong monodromy conjecture holds for $f$.
This can be reduced to [BaVe 26, Theorem C] using Proposition 3.3 below, since the strong monodromy conjecture for homogeneous polynomials with isolated singularities is easy. This proposition seems to hold also in the non-reduced case.
1. Topological zeta function
Let $\pi\,{:}\,\widetilde{Y}\,{\to}\,Y\,{:=}\,{\mathbb{C}}^{n}$ be an embedded resolution of singularities of $D\,{:=}\,f^{-1}(0)$ with $D_{i}$ the irreducible components of $\widetilde{D}\,{:=}\,\pi^{-1}(D)$. Let $N_{i}$, $\nu_{i}\thinspace{-}\thinspace 1$ be the multiplicities of $\pi^{*}f$ and $\pi^{*}{\rm d}x_{1}{\wedge}\cdots{\wedge}{\rm d}x_{n}$ respectively. Set $D_{I}\,{:=}\,\hbox{$\bigcap$}_{i\in I}\,D_{i}$, $D^{\circ}_{I}\,{:=}\,D_{I}\,{\setminus}\,\hbox{$\bigcup$}_{i\notin I}R_{i}$. The local topological zeta function is defined by
$Z^{\rm top}_{f,0}(s):=\hbox{$\sum$}_{I}\,\chi\bigl(D^{\circ}_{I}\cap\pi^{-1}(0)\bigr)\,\hbox{$\prod$}_{i\in I}\,(N_{i}s\thinspace{+}\thinspace\nu_{i})^{-1}.$
This is independent of a resolution, see [DeLo 92].
Remark 1.1.
In the curve case with n= 2n\,{=}\,2, the Euler number $\chi\bigl(D^{\circ}_{I}\cap\pi^{-1}(0)\bigr)$ vanishes if |I|= 1|I|\,{=}\,1 and $|\hbox{$\bigcup$}_{j\neq i}D_{i}\cap D_{j}|\,{=}\,2$, but the summands $\hbox{$\prod$}_{i\in I}\,(N_{i}s\thinspace{+}\thinspace\nu_{i})^{-1}$ for |I|= 2|I|\,{=}\,2 remain if $\hbox{$\bigcap$}_{i\in I}\,D_{i}\neq\emptyset$ and they may cancel each other out by using
$(s\thinspace{+}\thinspace\alpha)^{-1}\thinspace{-}\thinspace(s\thinspace{+}\thinspace\beta)^{-1}\,{=}\,(\beta\thinspace{-}\thinspace\alpha)(s\thinspace{+}\thinspace\alpha)^{-1}(s\thinspace{+}\thinspace\beta)^{-1}.$
In the reduced curve case it turns out that the poles of $Z^{\rm top}_{f,0}(s)$ are restricted to either 1-1 or $\nu_{i}/N_{i}$ for exceptional divisors $D_{i}\,{\subset}\,\pi^{-1}(0)$ with $|\hbox{$\bigcup$}_{j\neq i}D_{i}\cap D_{j}|\,{\geqslant}\,3$, see [Ve 95, Theorem 4.3].
2. Newton-nondegenerate case
We denote by $\Gamma_{\!+}(f)$ the Newton polytope  of $f$ at 0, which is the convex hull of the union of $\nu\thinspace{+}\thinspace{\mathbb{R}}_{\geqslant 0}^{n}$ for $\nu\,{\in}\,{\rm Supp}_{\bf x}\,f$, where
${\rm Supp}_{\bf x}\,f:=\{\nu\,{\in}\,{\mathbb{N}}^{n}\mid c_{\nu}\neq 0\thinspace\}\quad\hbox{for}\quad f\,{=}\,\hbox{$\sum$}_{\nu}\,c_{\nu}{\bf x}^{\nu}\,{\in}\,{\mathbb{C}}[{\bf x}],$
with ${\bf x}\,{=}\,(x_{1},\dots,x_{n})$ the coordinate system of ${\mathbb{C}}^{n}$. We say that $f$ is Newton nondegenerate, if for any compact  face $\sigma\,{\subset}\,\Gamma_{\!+}(f)$, we have
$\hbox{$\bigcap$}_{i=1}^{n}\,\bigl\{\thinspace x_{i}\partial_{x_{i}}f_{\sigma}\,{=}\,0\bigr\}\,{\subset}\,\{\thinspace x_{1}\cdots x_{n}\,{=}\,0\thinspace\},$
where $f_{\sigma}\,{:=}\,\hbox{$\bigoplus$}_{\nu\in\sigma}\,c_{\nu}{\bf x}^{\nu}$ with $c_{\nu}$ as above. We denote by $\Gamma_{\!f}$ the union of compact faces of $\Gamma_{\!+}(f)$. Set $d_{\sigma}\,{:=}\,\dim\sigma$ for $\sigma<\Gamma_{\!+}(f)$ (that is, $\sigma$ is a face of $\Gamma_{\!+}(f)$). We have the following.
Theorem 2.1 ([DeLo 92, Theorem 5.3]).
If $f$ is Newton-nondegenerate, there is the equality
(2.1) $Z^{\rm top}_{f,0}=\hbox{$\sum$}_{d_{\sigma}=0}\,J_{\sigma}(s)+\tfrac{s}{s+1}\,\hbox{$\sum$}_{d_{\sigma}>0}\,(-1)^{d_{\sigma}}d_{\sigma}!\thinspace{\rm Vol}(\sigma)J_{\sigma}(s),$
where the $\sigma$ are compact faces of $\Gamma_{\!+}(f)$, that is, $\sigma<\Gamma_{\!f}$.
We explain the notation only in the case n= 3n\,{=}\,3, see [DeLo 92], [LeVP 11] for the general case. (This is similar and easier when n= 2n\,{=}\,2.) Firstly $d_{\sigma}!\thinspace{\rm Vol}(\sigma)$ is the normalized volume  of $\sigma$. In the case $d_{\sigma}\,{=}\,0$ or 1, it is equal to 1 or $|\sigma\cap{\mathbb{N}}^{3}|\thinspace{-}\thinspace 1$ respectively. If $d_{\sigma}\,{=}\,2$ and $\sigma$ is simplicial so that the vertices of $\sigma$ are three points $a_{1},a_{2},a_{3}\,{\in}\,{\mathbb{N}}^{3}$, it coincides with the absolute value of the determinant of the matrix $(a_{1},a_{2},a_{3})$ divided by the lattice-theoretic distance  $N_{\sigma}$ between $\sigma$ and 0. The latter is defined by $N_{\sigma}\,{:=}\,\ell_{\sigma}(a_{i})$ for $i\,{\in}\,\{1,2,3\}$ with $\ell_{\sigma}$ the linear function with integral coefficients whose greatest common divisor is 1 and such that $\ell_{\sigma}$ is orthogonal to $\sigma$, that is, $\sigma\,{\subset}\,\ell_{\sigma}^{-1}(N_{\sigma})$. We put $\nu_{\sigma}\,{:=}\,\ell_{\sigma}(1,1,1)$, which is equal to the sum of the coefficients of $\ell_{\sigma}$. In general the normalized volume is defined by taking a partition of $\sigma$. Note that $\ell_{\sigma}$, $N_{\sigma}$, $\nu_{\sigma}$ are well defined for non-simplicial 2-dimensional $\sigma$. These can be defined also for non-compact faces (after taking the projection to ${\mathbb{R}}^{2}$ or $\mathbb{R}$} if necessary), where $N_{\sigma}\,{=}\,0$ or $1$ if two of the coefficients of $\ell_{\sigma}$ vanishes (since $f$ is reduced).
If $d_{\sigma}\,{=}\,2$, put
(2.2) $J_{\sigma}(s):=1/(N_{\sigma}s\thinspace{+}\thinspace\nu_{\sigma}).$
If $d_{\sigma}\,{=}\,1$ and $\sigma\,{=}\,\sigma_{1}\cap\sigma_{2}$ with $d_{\sigma_{i}}\,{=}\,2$ for i= 1,2i\,{=}\,1,2, set
(2.3) $J_{\sigma}(s):={\rm mult}(\ell_{\sigma_{1}},\ell_{\sigma_{2}})\big/\bigl((N_{\sigma_{1}}s\thinspace{+}\thinspace\nu_{\sigma_{1}})(N_{\sigma_{2}}s\thinspace{+}\thinspace\nu_{\sigma_{2}})\bigr).$
Here ${\rm mult}(\ell_{\sigma_{1}},\ell_{\sigma_{2}})$ is defined by the greatest common divisor of the absolute values of three minor determinants of size $2{\times}2$ of the matrix of size $3{\times}2$ defined by the coefficients of $\ell_{\sigma_{1}},\ell_{\sigma_{2}}$ (which can be identified with the exterior product $\ell_{\sigma_{1}}{\wedge}\,\ell_{\sigma_{2}}$).
If $d_{\sigma}\,{=}\,0$ and the number of 2-dimensional faces containing $\sigma$ is three, set
(2.4) $J_{\sigma}(s):={\rm mult}(\ell_{\sigma_{1}},\ell_{\sigma_{2}},\ell_{\sigma_{3}})\big/\bigl((N_{\sigma_{1}}s\thinspace{+}\thinspace\nu_{\sigma_{1}})(N_{\sigma_{2}}s\thinspace{+}\thinspace\nu_{\sigma_{2}})(N_{\sigma_{3}}s\thinspace{+}\thinspace\nu_{\sigma_{3}})\bigr).$
Here $\sigma_{1},\sigma_{2},\sigma_{3}$ are the three 2-dimensional faces containing $\sigma$ and ${\rm mult}(\ell_{\sigma_{1}},\ell_{\sigma_{2}},\ell_{\sigma_{3}})$ is the absolute value of the matrix defined by the coefficients of $\ell_{\sigma_{1}},\ell_{\sigma_{2}},\ell_{\sigma_{3}}$.
In general, let $\sigma_{i}$ for $i\in[1,r]$ be the 2-dimensional faces of $\Gamma_{\!+}(f)$ containing $\sigma$ and such that the $\sigma_{i}\cap\sigma_{i+1}$ for $i\in[1,r]$ are 1-dimensional faces of $\Gamma_{\!+}(f)$ with $\sigma_{r+1}\,{:=}\,\sigma_{r}$. Then
(2.5) $J_{\sigma}(s):=\hbox{$\sum$}_{i=3}^{r}\,{\rm mult}(\ell_{\sigma_{1}},\ell_{\sigma_{i-1}},\ell_{\sigma_{i}})\big/\bigl((N_{\sigma_{1}}s\thinspace{+}\thinspace\nu_{\sigma_{1}})(N_{\sigma_{i-1}}s\thinspace{+}\thinspace\nu_{\sigma_{i-1}})(N_{\sigma_{i}}s\thinspace{+}\thinspace\nu_{\sigma_{i}})\bigr).$
This can be defined more generally by choosing a simplicial partition of the cone spanned by $\ell_{\sigma_{1}},\dots,\ell_{\sigma_{r}}$. It is quite amazing that the sum is independent of the choice of a partition, see [DeLo 92, Lemma 5.1.1]. (This can be verified by using a computer in the situation of the proof of Theorem 3, see Remark 3.2 below.)
3. Proof of Theorem 3
Let $P,Q$ be the vertices of the unique compact face $\sigma_{0}$ of $\Gamma_{\!+}(f)$, which is 1-dimensional. Taking a monomial $x^{i}y^{j}z^{k}$ of highest degree which divides $f$, we see that their coordinates are given by
$P=(i\thinspace{+}\thinspace ar,j\thinspace{+}\thinspace br,k)\quad Q=(i,j,k\thinspace{+}\thinspace cr),$
where $a,b,c,r,i,j,k$ are integers such that $a,b,c,r$ are strictly positive, $c\,{=}\,a\thinspace{+}\thinspace b$, $a,b$ are coprime, and $i,j,k$ are 0 or 1, since $C$ is reduced.
The singular points of $C$ are [0: 1: 0][0\,{:}\,1\,{:}\,0] and [1: 0: 0][1\,{:}\,0\,{:}\,0], and the irreducible components of $C$ are rational curves passing through the two points, where the number of irreducible components is $r$. This may be verified by considering the factorization of a polynomial of one variable of degree $r$ whose coefficients are those of $f$. Since this polynomial must be reduced (otherwise $f$ is non-reduced), its roots are mutually distinct, hence $f$ is Newton-nondegenerate.
There are five 2-dimensional faces $\sigma_{i}$ of $\Gamma_{\!+}(f)$ with $i\,{\in}\,[1,5]$. The coefficients of their corresponding linear functions $\ell^{(i)}$ for $i\,{\in}\,[1,5]$ are respectively
$(1,0,0),\quad(0,1,0),\quad(0,0,1),\quad(c,0,a),\quad(0,c,b),$
and the corresponding $(N_{\sigma_{i}},\nu_{\sigma_{i}})$, which will be denoted by $(N_{i},\nu_{i})$, are respectively
$(i,1),\quad(j,1),\quad(k,1),\quad(m,a\thinspace{+}\thinspace c),\quad(n,b\thinspace{+}\thinspace c),$
$\hbox{with}\quad\quad m\,{:=}\,ci\thinspace{+}\thinspace ak\thinspace{+}\thinspace acr,\quad n\,{:=}\,cj\thinspace{+}\thinspace bk\thinspace{+}\thinspace bcr.$
We then see that $J_{\sigma_{0}}(s)$, $J_{P}(s)$, $J_{Q}(s)$ are given by
$J_{\sigma_{0}}(s)$ $=c\big/\bigl((ms\thinspace{+}\thinspace a\thinspace{+}\thinspace c)(ns\thinspace{+}\thinspace b\thinspace{+}\thinspace c)\bigr),$
$J_{P}(s)$ $=c^{2}\big/\bigl((ks\thinspace{+}\thinspace 1)(ns\thinspace{+}\thinspace b\thinspace{+}\thinspace c)(ms\thinspace{+}\thinspace a\thinspace{+}\thinspace c)\bigr),$
$J_{Q}(s)$ $=b\big/\bigl((is\thinspace{+}\thinspace 1)(js\thinspace{+}\thinspace 1)(ns\thinspace{+}\thinspace b\thinspace{+}\thinspace c)\bigr)$
$\quad+ac\big/\bigl((ns\thinspace{+}\thinspace b\thinspace{+}\thinspace c)(ms\thinspace{+}\thinspace a\thinspace{+}\thinspace c)(is\thinspace{+}\thinspace 1)\bigr).$
Indeed, for the first two equalities it is enough to calculate the minor determinants of $\bigl(\begin{smallmatrix}c&0&a\\ 0&c&b\end{smallmatrix}\bigr)$ and the determinant of $\bigl(\begin{smallmatrix}0&0&1\\ c&0&a\\ 0&c&b\end{smallmatrix}\bigr)$. For the last one, we verify that the 2-dimensional faces containing $Q$ are $\sigma_{1}$, $\sigma_{2}$, $\sigma_{5}$, $\sigma_{4}$, and compute the determinants of $\bigl(\begin{smallmatrix}1&0&0\\ 0&1&0\\ 0&c&b\end{smallmatrix}\bigr)$ and $\bigl(\begin{smallmatrix}0&c&b\\ c&0&a\\ 1&0&0\end{smallmatrix}\bigr)$. (Recall that there is another expression, but these coincide according to a computer algebra program, see Remark 3.2 below.)
It remains to show that the poles $(a\thinspace{+}\thinspace c)/m$, $(b\thinspace{+}\thinspace c)/n$ are poles of the topological zeta function of the singular points of $C$, but this is clear by the formula (2.1) for the curve case since $\ell_{4}$, $\ell_{5}$ are pullbacks of the linear functions for the image of $\sigma_{0}$ by projections to ${\mathbb{R}}^{2}$. This completes the proof of Theorem 3.
Remark 3.1.
In the case $f$ is a homogeneous polynomial, an embedded resolution can be obtained by the blowup of the origin followed by the base change of an embedded resolution of the projective hypersurface by the projection to the exceptional divisor of the first blowup. In view of the definition of local topological zeta function, this implies that 3/d-3/d is the only possible pole of $Z^{\rm top}_{f,0}(s)$ which is not a pole of $Z^{\rm top}_{h,0}(s)$ for a local defining function $h$ of $C$.
Remark 3.2.
The above argument can be used to prove the formula (2.1) for our case by comparing it with the one for the two variable case. What is very amazing is that the numerator  of the (modified) “global” topological zeta functions of $C$ (which contains the contribution from ${\mathbb{P}}^{2}\,{\setminus}\,C$) is divisible  by $s+3/d$ so that the pole 3/d-3/d disappears. Note that we need the division by $ds\thinspace{+}\thinspace 3$ for the passage from the (modified) “global” zeta functions of $C$ to $Z^{\rm top}_{f,0}(s)$ (via the blowup at 0) by definition. Using Macaulay2 for instance, this can be seen as follows.
S=QQ[a,b,i,j,k,r,s]; c=a+b; m=c*i+c*a*r+a*k; n=c*j+c*b*r+b*k; d=i+j+k+c*r;
Z = -s/(s+1) *c*r/(m*s+a+c)/(n*s+b+c) + b/(i*s+1)/(j*s+1)/(n*s+b+c) +
a*c/(n*s+b+c)/(m*s+a+c)/(i*s+1) + c^2/(k*s+1)/(n*s+b+c)/(m*s+a+c);
Z1 = -s/(s+1)*r/(m*s+a+c) + c/(m*s+a+c)/(k*s+1) + a/(m*s+a+c)/(i*s+1);
Z2 = -s/(s+1)*r/(n*s+b+c) + c/(n*s+b+c)/(k*s+1) + b/(n*s+b+c)/(j*s+1);
R = 1/(i*s+1)/(j*s+1); Z - (Z1+Z2+R)/(d*s+3)
Here Z, Z1, Z2 are the local topological zeta functions of $f$ and the two singular points of $C$, and R is the “remaining” part of the (modified) “global” zeta function of $C$. (Notice that R is independent of $k\,{\in}\,\{0,1\}$, since $\{z\,{=}\,0\}\,{\setminus}\,{\rm Sing}\,C\cong{\mathbb{C}}^{*}$. In the cases (i,j)=(0,0)(i,j)\,{=}\,(0,0), $(1,0)$, and $(1,1)$, R is the contribution from ${\mathbb{P}}^{2}\,{\setminus}\,C$, $\{x\,{=}\,0\}\,{\setminus}\,{\rm Sing}\,C\,({\cong}\,{\mathbb{C}})$, and [0: 0: 1][0\,{:}\,0\,{:}\,1] respectively.) This computation shows that Z1+Z2+R is divisible by d*s+3, and the quotient coincides with Z in the field of fractions of the polynomial ring. This gives also a “proof” of (2.1) for our case assuming it in the two variable case (and using a computer algebra program).
It is not recommended to remove ; at the end of the definition of Z. If one prefers, S=QQ[a,b,i,j,k,r,s]; may be replaced by S=QQ[s]; a=2; b=3; i=0; j=0; k=0; r=1; for instance. The coincidence with another expression of $J_{Q}(s)$ can be seen by the following:
b/(i*s+1)/(j*s+1)/(n*s+b+c) + a*c/(n*s+b+c)/(m*s+a+c)/(i*s+1) -
(b*c/(j*s+1)/(n*s+b+c)/(m*s+a+c) + a/(m*s+a+c)/(i*s+1)/(j*s+1))
We note here the following (which is needed for a proof of Proposition 1).
Proposition 3.3.
Assume a reduced projective hypersurface $Z$ of dimension at least $2$ and degree at least $3$ has only homogeneous isolated singularities which are not ordinary double points, and there is a nonzero linear function $\ell(\nu_{1},\dots,\nu_{n})\,{=}\,\hbox{$\sum$}_{i=1}^{n}\,c_{i}\nu_{i}$ on ${\mathbb{R}}^{n}$ with rational coefficients $c_{i}$ such that ${\rm Supp}_{\bf x}\thinspace f\,{\subset}\,\ell^{-1}(0)$ with $f$ a defining polynomial of $Z$. Then $Z$ is a cone.
Proof.
[…]
We thus get that the $c_{j}$ are independent of $j\,{\neq}\,i$. These must vanish, since $f^{(j)}$ has non-isolated singularities otherwise. Then only $c_{i}$ is nonzero, that is, $f$ is a polynomial of $x_{j}$ for $j\,{\neq}\,i$. This contradicts the assumption that $Z$ is not a cone. So the assertion follows. ∎