### v4

# Strong monodromy conjecture for defining polynomials of projective hypersurfaces having only weighted homogeneous isolated singularities

Abstract.
Let $Z\subset{\bf P}^{n-1}$ be a hypersurface such that the associated reduced hypersurface $Z_{\rm red}$ has only weighted homogeneous isolated singularities. In the case $Z$ is a reduced curve or $Z_{\rm red}$ has only homogeneous isolated singularities with $n$ at least $4$, we show that the strong monodromy conjecture for a defining polynomial $f$ of $Z$ follows immediately from arxiv:1609.04801v1 using in the reduced curve case a formula of Denef and Loeser for Newton-nondegenerate polynomials of three variables (which can be deduced in the applied case from the one for the two variable case) together with known results about the strong monodromy conjecture in the two variable case. Here an amazing cancellation occurs so that possible counterexamples fail.
Introduction
Let $f$ be a defining polynomial of a projective hypersurface $Z\,{\subset}\,{\mathbb{P}}^{n-1}$ such that the associated reduced hypersurface $Z_{\rm red}$ has only weighted homogeneous singularities. Set $d\,{:=}\,\deg Z$. We have the following (which is an immediate consequence of [Sa 16b, Theorem 2] and [Sa 07, Theorem 2], see also [BaVe 26, Theorem C] including the non-reduced case).
Theorem 1.
Assume $Z$ is reduced and there is not a nonzero vector field of degree 0 which annihilates $f$ and whose trace is nonzero. Then $\tfrac{n}{d}$ is a root of the Bernstein-Sato polynomial $b_{f}(s)$ of $f$.
Here the trace of a vector field of degree 0 means the trace of the matrix consisting of the coefficients $c_{i,j}$ of a vector field $\hbox{$\sum$}_{i,j}\,c_{i,j}x_{i}\partial_{x_{j}}$ of degree 0 with $c_{i,j}\,{\in}\,{\mathbb{C}}$. This theorem immediately follows from the $E_{2}$-degeneration of the pole order spectral sequence (see [Sa 16b, Theorem 2] using [Sa 07, Theorem 2], since the image of the $E_{1}$-differential ${\rm d}_{1}$ is given by the traces of annihilating vector fields of degree 0, see Remark 3.5 below.
Employing only Jordan normal form, we can verify in a completely elementary way the following (which is a natural generalization of an argument in the proof of [Sa 25, Proposition 8.1], see also [BaVe 26, Lemma 3.11]).
Lemma 1.
Assume $Z$ is reduced. If $f$ is annihilated by a nonzero vector field $\xi$ of degree 0, then $f$ is also annihilated by the semisimple part $\xi^{\prime}$ of $\xi$ ((which is defined by using Jordan normal form)).
The trace of a vector field of degree 0 is zero if its semisimple part vanishes. So Lemma 1 implies the following (which can replace [Sa 25, Propositions 8.1-2]).
Corollary 1.
Assume $Z$ is reduced. If the second hypothesis of Theorem 1 is not satisfied, then $f$ is extremely degenerated, that is, it is annihilated by a nonzero diagonal vector field of degree 0 of the form $\hbox{$\sum$}_{i=0}^{n}\,c_{i}x_{i}\partial_{i}$ with $c_{i}\,{\in}\,{\mathbb{Q}}$.
Here we have $c_{i}\,{\in}\,{\mathbb{Q}}$ by using a $\mathbb{Q}$}-basis $\gamma_{j}$ of $\hbox{$\sum$}_{i=0}^{n}\,{\mathbb{Q}}\,c_{i}\,{\subset}\,{\mathbb{C}}$ so that $c_{i}\,{=}\,\hbox{$\sum$}_{j}\,\gamma_{j}c^{\prime}_{j,i}$, where we may get several annihilating vector fields of degree 0 with rational coefficients. (Note that monomials are $\mathbb{C}$}-linearly independent  in the polynomial ring. This is the key point in the proof of Lemma 1.)
In the reduced curve case, using Theorem 1 and Corollary 1, it is easy to reduce the strong monodromy conjecture  for the local topological zeta function  $Z^{\rm top}_{f,0}(s)$ (claiming that a pole of $Z^{\rm top}_{f,0}(s)$ is a root of the Bernstein-Sato polynomial $b_{f}(s)$, see [DeLo 92], [BSY 11], [Sa 16a], [Wa 17], [BaVe 26] among others) to the following (using also some known results in the two variable case [Lo 88], see Remark 3.1 below).
Theorem 2.
If $Z$ is a reduced curve $C$ and is extremely degenerated, then any pole of the local topological zeta function $Z^{\rm top}_{f,0}(s)$ is a pole of the local topological zeta function of a singular point of $C$.
This can be verified rather easily if one admits a formula for the local topological zeta function of a Newton-nondegenerate polynomial of three variables, see [DeLo 92] (which is quite amazing). In the applied case it is possible to deduce this formula from the one for the two variable case (which seems much easier to prove) using the blowup at the origin followed by the ones with line centers together with some computations of rational functions by a computer algebra program such as Macaulay2 or Singular among others. Here it is very surprising that a pole $\tfrac{1}{ds+3}$ disappears  (more precisely, the numerator is divisible by $ds\thinspace{+}\thinspace 3$), when one calculates $Z^{\rm top}_{f,0}(s)$ using a desingularization even if the Euler number of ${\mathbb{P}}^{2}\,{\setminus}\,C$ does not vanish so that one could expect a counterexample, see Remark 3.2 below. Note also that a solution of the strong monodromy conjecture for the local topological zeta function in a reduced case implies that for a non-reduced case if the multiplicities are constant, see Remark 1 below.
The strong monodromy conjecture has been shown quite recently in the not necessarily reduced curve case under the hypothesis that the associated reduced curve has only weighted homogeneous singularities (see [BaVe 26]), where the vanishing of the residue at $\tfrac{3}{d}$ is proved in the needed case. Here it does not seem to be mentioned that the following assertion actually holds.
Theorem 3.
Let $Z\,{\subset}\,{\mathbb{P}}^{n-1}$ be a projective hypersurface defined by a polynomial $f$ with $n\,{\geqslant}\,4$. Assume the reduced hypersurface $Z_{\rm red}$ has only homogeneous isolated singularities. Then the strong monodromy conjecture holds for the local topological zeta function of $f$.
This is reduced to Theorem 1 and Corollary 1 using Proposition 3.3 and Remark 1 below, since the strong monodromy conjecture for a homogeneous polynomial with an isolated singularity is easy. Note that $Z$ is irreducible since $n\,{\geqslant}\,4$.
Remark 1.
For a positive integer $m$, the poles of the local topological zeta function for $f^{m}$ are those for $f$ divided by $m$. Indeed, the $N_{i}$ are multiplied by $m$ (although the $\nu_{i}$ do not change) by replacing $f$ with $f^{m}$ in the notation of Section 1 below. Moreover the roots of the Bernstein-Sato polynomial $b_{f}(s)$ divided by $m$ are contained in the roots of $b_{f^{m}}(s)$, see the last divisibility in Remark 3.4 below. So the strong monodromy conjecture for the local topological zeta function of $f^{m}$ is reduced to that for $f$.
We thank S.-J. Jung, D. Bath, and N. Budur for useful comments about this note.
1. Topological zeta function
Let $\pi\,{:}\,\widetilde{Y}\,{\to}\,Y\,{:=}\,{\mathbb{C}}^{n}$ be an embedded resolution of singularities of $D\,{:=}\,f^{-1}(0)$ with $D_{i}$ the irreducible components of $\widetilde{D}\,{:=}\,\pi^{-1}(D)$. Let $N_{i}$, $\nu_{i}\thinspace{-}\thinspace 1$ be the multiplicities of $\pi^{*}f$ and $\pi^{*}{\rm d}x_{1}{\wedge}\cdots{\wedge}{\rm d}x_{n}$ respectively. Set $D_{I}\,{:=}\,\hbox{$\bigcap$}_{i\in I}\,D_{i}$, $D^{\circ}_{I}\,{:=}\,D_{I}\,{\setminus}\,\hbox{$\bigcup$}_{i\notin I}R_{i}$. The local topological zeta function is defined by
$Z^{\rm top}_{f,0}(s):=\hbox{$\sum$}_{I}\,\chi\bigl(D^{\circ}_{I}\cap\pi^{-1}(0)\bigr)\,\hbox{$\prod$}_{i\in I}\,(N_{i}s\thinspace{+}\thinspace\nu_{i})^{-1}.$
This is independent of a resolution, see [DeLo 92].
Remark 1.1.
In the curve case with n= 2n\,{=}\,2, the Euler number $\chi\bigl(D^{\circ}_{I}\cap\pi^{-1}(0)\bigr)$ vanishes if |I|= 1|I|\,{=}\,1 and $|\hbox{$\bigcup$}_{j\neq i}D_{i}\cap D_{j}|\,{=}\,2$, but the summands $\hbox{$\prod$}_{i\in I}\,(N_{i}s\thinspace{+}\thinspace\nu_{i})^{-1}$ for |I|= 2|I|\,{=}\,2 remain if $\hbox{$\bigcap$}_{i\in I}\,D_{i}\neq\emptyset$ and they may cancel each other out by using
$(s\thinspace{+}\thinspace\alpha)^{-1}\thinspace{-}\thinspace(s\thinspace{+}\thinspace\beta)^{-1}\,{=}\,(\beta\thinspace{-}\thinspace\alpha)(s\thinspace{+}\thinspace\alpha)^{-1}(s\thinspace{+}\thinspace\beta)^{-1}.$
In the reduced curve case it turns out that the poles of $Z^{\rm top}_{f,0}(s)$ are restricted to either 1-1 or $\nu_{i}/N_{i}$ for exceptional divisors $D_{i}\,{\subset}\,\pi^{-1}(0)$ with $|\hbox{$\bigcup$}_{j\neq i}D_{i}\cap D_{j}|\,{\geqslant}\,3$, see [Ve 95, Theorem 4.3].
2. Newton-nondegenerate case
We denote by $\Gamma_{\!+}(f)$ the Newton polytope  of $f$ at 0, which is the convex hull of the union of $\nu\thinspace{+}\thinspace{\mathbb{R}}_{\geqslant 0}^{n}$ for $\nu\,{\in}\,{\rm Supp}_{\bf x}\,f$, where
${\rm Supp}_{\bf x}\,f:=\{\nu\,{\in}\,{\mathbb{N}}^{n}\mid c_{\nu}\neq 0\thinspace\}\quad\hbox{for}\quad f\,{=}\,\hbox{$\sum$}_{\nu}\,c_{\nu}{\bf x}^{\nu}\,{\in}\,{\mathbb{C}}[{\bf x}],$
with ${\bf x}\,{=}\,(x_{1},\dots,x_{n})$ the coordinate system of ${\mathbb{C}}^{n}$. We say that $f$ is Newton nondegenerate, if for any compact  face $\sigma\,{\subset}\,\Gamma_{\!+}(f)$, we have
$\hbox{$\bigcap$}_{i=1}^{n}\,\bigl\{\thinspace x_{i}\partial_{x_{i}}f_{\sigma}\,{=}\,0\bigr\}\,{\subset}\,\{\thinspace x_{1}\cdots x_{n}\,{=}\,0\thinspace\},$
where $f_{\sigma}\,{:=}\,\hbox{$\bigoplus$}_{\nu\in\sigma}\,c_{\nu}{\bf x}^{\nu}$ with $c_{\nu}$ as above. We denote by $\Gamma_{\!f}$ the union of compact faces of $\Gamma_{\!+}(f)$. Set $d_{\sigma}\,{:=}\,\dim\sigma$ for $\sigma<\Gamma_{\!+}(f)$ (that is, $\sigma$ is a face of $\Gamma_{\!+}(f)$). We have the following.
Theorem 2.1 ([DeLo 92, Theorem 5.3]).
If $f$ is Newton-nondegenerate, there is the equality
(2.1) $Z^{\rm top}_{f,0}=\hbox{$\sum$}_{d_{\sigma}=0}\,J_{\sigma}(s)+\tfrac{s}{s+1}\,\hbox{$\sum$}_{d_{\sigma}>0}\,(-1)^{d_{\sigma}}d_{\sigma}!\thinspace{\rm Vol}(\sigma)J_{\sigma}(s),$
where the $\sigma$ are compact faces of $\Gamma_{\!+}(f)$, that is, $\sigma<\Gamma_{\!f}$.
We explain the notation in the case n= 3n\,{=}\,3, see [DeLo 92], [LeVP 11] for the general case. (This is similar and easier when n= 2n\,{=}\,2.) Firstly $d_{\sigma}!\thinspace{\rm Vol}(\sigma)$ is the normalized volume  of $\sigma$. In the case $d_{\sigma}\,{=}\,0$ or 1, it is equal to 1 or $|\sigma\cap{\mathbb{N}}^{3}|\thinspace{-}\thinspace 1$ respectively. If $d_{\sigma}\,{=}\,2$ and $\sigma$ is simplicial so that the vertices of $\sigma$ are three points $a_{1},a_{2},a_{3}\,{\in}\,{\mathbb{N}}^{3}$, it coincides with the absolute value of the determinant of the matrix $(a_{1},a_{2},a_{3})$ divided by the lattice-theoretic distance  $N_{\sigma}$ between $\sigma$ and 0. The latter is defined by $N_{\sigma}\,{:=}\,\ell_{\sigma}(a_{i})$ for $i\,{\in}\,\{1,2,3\}$ with $\ell_{\sigma}$ the linear function with integral coefficients whose greatest common divisor is 1 and such that $\ell_{\sigma}$ is orthogonal to $\sigma$, that is, $\sigma\,{\subset}\,\ell_{\sigma}^{-1}(N_{\sigma})$. We put $\nu_{\sigma}\,{:=}\,\ell_{\sigma}(1,1,1)$, which is equal to the sum of the coefficients of $\ell_{\sigma}$. In general the normalized volume is defined by taking a partition of $\sigma$. Note that $\ell_{\sigma}$, $N_{\sigma}$, $\nu_{\sigma}$ are well defined for non-simplicial 2-dimensional $\sigma$. These can be defined also for non-compact faces (after taking the projection to ${\mathbb{R}}^{2}$ or $\mathbb{R}$} if necessary), where $N_{\sigma}\,{=}\,0$ or $1$ if two of the coefficients of $\ell_{\sigma}$ vanishes (since $f$ is reduced).
If $d_{\sigma}\,{=}\,2$, put
$J_{\sigma}(s):=1/(N_{\sigma}s\thinspace{+}\thinspace\nu_{\sigma}).$
If $d_{\sigma}\,{=}\,1$ and $\sigma\,{=}\,\sigma_{1}\cap\sigma_{2}$ with $d_{\sigma_{i}}\,{=}\,2$ for i= 1,2i\,{=}\,1,2, set
$J_{\sigma}(s):={\rm mult}(\ell_{\sigma_{1}},\ell_{\sigma_{2}})\big/\bigl((N_{\sigma_{1}}s\thinspace{+}\thinspace\nu_{\sigma_{1}})(N_{\sigma_{2}}s\thinspace{+}\thinspace\nu_{\sigma_{2}})\bigr).$
Here ${\rm mult}(\ell_{\sigma_{1}},\ell_{\sigma_{2}})$ is defined by the greatest common divisor of the absolute values of three minor determinants of size $2{\times}2$ of the matrix of size $3{\times}2$ defined by the coefficients of $\ell_{\sigma_{1}},\ell_{\sigma_{2}}$ (which can be identified with the exterior product $\ell_{\sigma_{1}}{\wedge}\,\ell_{\sigma_{2}}$).
If $d_{\sigma}\,{=}\,0$ and the number of 2-dimensional faces containing $\sigma$ is three, set
$J_{\sigma}(s):={\rm mult}(\ell_{\sigma_{1}},\ell_{\sigma_{2}},\ell_{\sigma_{3}})\big/\bigl((N_{\sigma_{1}}s\thinspace{+}\thinspace\nu_{\sigma_{1}})(N_{\sigma_{2}}s\thinspace{+}\thinspace\nu_{\sigma_{2}})(N_{\sigma_{3}}s\thinspace{+}\thinspace\nu_{\sigma_{3}})\bigr).$
Here $\sigma_{1},\sigma_{2},\sigma_{3}$ are the three 2-dimensional faces containing $\sigma$ and ${\rm mult}(\ell_{\sigma_{1}},\ell_{\sigma_{2}},\ell_{\sigma_{3}})$ is the absolute value of the matrix defined by the coefficients of $\ell_{\sigma_{1}},\ell_{\sigma_{2}},\ell_{\sigma_{3}}$.
In general, let $\sigma_{i}$ for $i\in[1,r]$ be the 2-dimensional faces of $\Gamma_{\!+}(f)$ containing $\sigma$ and such that the $\sigma_{i}\cap\sigma_{i+1}$ for $i\in[1,r]$ are 1-dimensional faces of $\Gamma_{\!+}(f)$ with $\sigma_{r+1}\,{:=}\,\sigma_{r}$. Then
$J_{\sigma}(s):=\hbox{$\sum$}_{i=3}^{r}\,{\rm mult}(\ell_{\sigma_{1}},\ell_{\sigma_{i-1}},\ell_{\sigma_{i}})\big/\bigl((N_{\sigma_{1}}s\thinspace{+}\thinspace\nu_{\sigma_{1}})(N_{\sigma_{i-1}}s\thinspace{+}\thinspace\nu_{\sigma_{i-1}})(N_{\sigma_{i}}s\thinspace{+}\thinspace\nu_{\sigma_{i}})\bigr).$
This can be defined more generally by choosing a simplicial partition of the cone spanned by $\ell_{\sigma_{1}},\dots,\ell_{\sigma_{r}}$. It is quite amazing that the sum is independent of the choice of a partition, see [DeLo 92, Lemma 5.1.1]. (This can be verified by using a computer in the situation of the proof of Theorem 2, see Remark 3.2 below.)
3. Proofs of Theorem 2 and Lemma 1
Proof of Theorem 2.
Let $P,Q$ be the vertices of the unique compact face $\sigma_{0}$ of $\Gamma_{\!+}(f)$, which is 1-dimensional. Taking a monomial $x^{i}y^{j}z^{k}$ of highest degree which divides $f$, we see that their coordinates are given by
$P=(i\thinspace{+}\thinspace ar,j\thinspace{+}\thinspace br,k)\quad Q=(i,j,k\thinspace{+}\thinspace cr),$
where $a,b,c,r,i,j,k$ are integers such that $a,b,c,r$ are strictly positive, $c\,{=}\,a\thinspace{+}\thinspace b$, $a,b$ are coprime, and $i,j,k$ are 0 or 1, since $C$ is reduced.
The singular points of $C$ are [0: 1: 0][0\,{:}\,1\,{:}\,0] and [1: 0: 0][1\,{:}\,0\,{:}\,0], and the irreducible components of $C$ are rational curves passing through the two points, where the number of irreducible components is $r$. This may be verified by considering the factorization of a polynomial of one variable of degree $r$ whose coefficients are those of $f$. Since this polynomial must be reduced (otherwise $f$ is non-reduced), its roots are mutually distinct, hence $f$ is Newton-nondegenerate.
There are five 2-dimensional faces $\sigma_{i}$ of $\Gamma_{\!+}(f)$ with $i\,{\in}\,[1,5]$. The coefficients of their corresponding linear functions $\ell^{(i)}$ for $i\,{\in}\,[1,5]$ are respectively
$(1,0,0),\quad(0,1,0),\quad(0,0,1),\quad(c,0,a),\quad(0,c,b),$
and the corresponding $(N_{\sigma_{i}},\nu_{\sigma_{i}})$, which will be denoted by $(N_{i},\nu_{i})$, are respectively
$(i,1),\quad(j,1),\quad(k,1),\quad(m,a\thinspace{+}\thinspace c),\quad(n,b\thinspace{+}\thinspace c),$
$\hbox{with}\quad\quad m\,{:=}\,ci\thinspace{+}\thinspace ak\thinspace{+}\thinspace acr,\quad n\,{:=}\,cj\thinspace{+}\thinspace bk\thinspace{+}\thinspace bcr.$
We then see that $J_{\sigma_{0}}(s)$, $J_{P}(s)$, $J_{Q}(s)$ are given by
$J_{\sigma_{0}}(s)$ $=c\big/\bigl((ms\thinspace{+}\thinspace a\thinspace{+}\thinspace c)(ns\thinspace{+}\thinspace b\thinspace{+}\thinspace c)\bigr),$
$J_{P}(s)$ $=c^{2}\big/\bigl((ks\thinspace{+}\thinspace 1)(ns\thinspace{+}\thinspace b\thinspace{+}\thinspace c)(ms\thinspace{+}\thinspace a\thinspace{+}\thinspace c)\bigr),$
$J_{Q}(s)$ $=b\big/\bigl((is\thinspace{+}\thinspace 1)(js\thinspace{+}\thinspace 1)(ns\thinspace{+}\thinspace b\thinspace{+}\thinspace c)\bigr)$
$\quad+ac\big/\bigl((ns\thinspace{+}\thinspace b\thinspace{+}\thinspace c)(ms\thinspace{+}\thinspace a\thinspace{+}\thinspace c)(is\thinspace{+}\thinspace 1)\bigr).$
Indeed, for the first two equalities it is enough to calculate the minor determinants of $\bigl(\begin{smallmatrix}c&0&a\\ 0&c&b\end{smallmatrix}\bigr)$ and the determinant of $\bigl(\begin{smallmatrix}0&0&1\\ c&0&a\\ 0&c&b\end{smallmatrix}\bigr)$. For the last one, we verify that the 2-dimensional faces containing $Q$ are $\sigma_{1}$, $\sigma_{2}$, $\sigma_{5}$, $\sigma_{4}$, and compute the determinants of $\bigl(\begin{smallmatrix}1&0&0\\ 0&1&0\\ 0&c&b\end{smallmatrix}\bigr)$ and $\bigl(\begin{smallmatrix}0&c&b\\ c&0&a\\ 1&0&0\end{smallmatrix}\bigr)$. (Recall that there is another expression, but these coincide according to a computer algebra program, see Remark 3.2 below.)
It remains to show that the poles $(a\thinspace{+}\thinspace c)/m$, $(b\thinspace{+}\thinspace c)/n$ are poles of the topological zeta function of the singular points of $C$, but this is clear by the formula (2.1) for the curve case since $\ell_{4}$, $\ell_{5}$ are pullbacks of the linear functions for the image of $\sigma_{0}$ by projections to ${\mathbb{R}}^{2}$. This completes the proof of Theorem 2. ∎
Remark 3.1.
In the case $f$ is a homogeneous polynomial, an embedded resolution can be obtained by the blowup of the origin followed by the base change of an embedded resolution of the projective hypersurface by the projection to the exceptional divisor of the first blowup. In view of the definition of local topological zeta function, this implies that $\tfrac{3}{d}$ is the only possible pole of $Z^{\rm top}_{f,0}(s)$ which is not a pole of $Z^{\rm top}_{h,0}(s)$ for a local defining function $h$ of $C$.
Remark 3.2.
The above argument can be used to prove the formula (2.1) for our case by comparing it with the one for the two variable case. What is very amazing is that the numerator  of the (modified) “global” topological zeta functions of $C$ (which contains the contribution from ${\mathbb{P}}^{2}\,{\setminus}\,C$) is divisible  by $s+\tfrac{3}{d}$ so that the pole $\tfrac{3}{d}$ disappears. Note that we need the division by $ds\thinspace{+}\thinspace 3$ for the passage from the (modified) “global” zeta functions of $C$ to $Z^{\rm top}_{f,0}(s)$ (via the blowup at 0) by definition. Using Macaulay2 for instance, this can be seen as follows.
S=QQ[a,b,i,j,k,r,s]; c=a+b; m=c*i+c*a*r+a*k; n=c*j+c*b*r+b*k; d=i+j+k+c*r;
Z = -s/(s+1) *c*r/(m*s+a+c)/(n*s+b+c) + b/(i*s+1)/(j*s+1)/(n*s+b+c) +
a*c/(n*s+b+c)/(m*s+a+c)/(i*s+1) + c^2/(k*s+1)/(n*s+b+c)/(m*s+a+c);
Z1 = -s/(s+1)*r/(m*s+a+c) + c/(m*s+a+c)/(k*s+1) + a/(m*s+a+c)/(i*s+1);
Z2 = -s/(s+1)*r/(n*s+b+c) + c/(n*s+b+c)/(k*s+1) + b/(n*s+b+c)/(j*s+1);
R = 1/(i*s+1)/(j*s+1); Z - (Z1+Z2+R)/(d*s+3)
Here Z, Z1, Z2 are the local topological zeta functions of $f$ and the two singular points of $C$, and R is the “remaining” part of the (modified) “global” zeta function of $C$. Notice that R is independent of $k\,{\in}\,\{0,1\}$, since $\{z\,{=}\,0\}\,{\setminus}\,{\rm Sing}\,C\cong{\mathbb{C}}^{*}$. In the cases (i,j)=(0,0)(i,j)\,{=}\,(0,0), $(1,0)$, and $(1,1)$, R is the contribution from ${\mathbb{P}}^{2}\,{\setminus}\,C$, $\{x\,{=}\,0\}\,{\setminus}\,{\rm Sing}\,C\,({\cong}\,{\mathbb{C}})$, and [0: 0: 1][0\,{:}\,0\,{:}\,1] respectively. This computation shows that Z1+Z2+R is divisible by d*s+3, and the quotient coincides with Z in the field of fractions of the polynomial ring. This gives also a “proof” of (2.1) for our case assuming it in the two variable case (and using a computer algebra program).
It is not recommended to remove ; at the end of the definition of Z. If one prefers, S=QQ[a,b,i,j,k,r,s]; may be replaced by S=QQ[s]; a=2; b=3; i=0; j=0; k=0; r=1; for instance. The coincidence with another expression of $J_{Q}(s)$ can be seen by the following:
b/(i*s+1)/(j*s+1)/(n*s+b+c) + a*c/(n*s+b+c)/(m*s+a+c)/(i*s+1) -
(b*c/(j*s+1)/(n*s+b+c)/(m*s+a+c) + a/(m*s+a+c)/(i*s+1)/(j*s+1))
As for the relation to the trace of the logarithmic vector field in [BaVe 26], it may be seen by comparing the following:
u = -(3*m - d*(a+c))
v = 3*n - d*(b+c)
p = -a*(d-3*j) + b*(d-3*i)
q = (a*r*j-b*r*i + c*r*j+b*r*k+b*c*r^2 - (c*r*i+a*r*k+a*c*r^2))/r
All of these turn out to be the same polynomial. The last one is the sum of the coefficients of a linear function $\ell$ vanishing on $P$ and $Q$, where $\ell$ is identified with a vector field $\xi$ of degree 0 annihilating $f$.
We note here the following (which is needed for a proof of Theorem 1).
Proposition 3.3.
Let $Z$ be a reduced projective hypersurface of dimension at least $2$. Assume the hypersurface has only homogeneous isolated singularities and there is a nonzero linear function $\ell(\nu_{1},\dots,\nu_{n})\,{=}\,\hbox{$\sum$}_{i=1}^{n}\,c_{i}\nu_{i}$ on ${\mathbb{R}}^{n}$ with rational coefficients such that ${\rm Supp}_{\bf x}\thinspace f\,{\subset}\,\ell^{-1}(0)$ with $f$ a defining polynomial of $Z$. Then $Z$ is a cone or smooth with $d\,{:=}\,\deg f\,{=}\,2$.
Proof.
[…]
$\hbox{$\sum$}_{k=1}^{\mu}\,t^{\alpha_{k}}=\hbox{$\prod$}_{j\neq i}\,(t\thinspace{-}\thinspace t^{w_{j}})/(t^{w_{j}}\thinspace{-}\thinspace 1),\quad\mu=\hbox{$\prod$}_{j\neq i}\,(w_{j}^{-1}\thinspace{-}\thinspace 1),$
where the $w^{(i)}_{j}$ are denoted by $w_{j}$. We then see that the weights $w^{(i)}_{j}$ must be constant, since the weights in the homogeneous case with $\mu\,{\neq}\,1$ can be read from the $\alpha_{k}\thinspace{-}\thinspace\alpha_{1}$ for $k\,{\in}\,[2,n]$.
We thus proved that the $c_{j}$ are independent of $j\,{\neq}\,i$. This implies the inclusion
${\rm Supp}_{\bf x}\thinspace f\subset\bigl\{\nu\,{\in}\,{\mathbb{N}}^{n}\mid\hbox{$\sum$}_{j}\,\nu_{j}\,{=}\,d,\,\nu_{i}\,{=}\,c\thinspace\hbox{$\sum$}_{j\,{\neq}\,i}\,\nu_{j}\bigr\}$
for some $c\,{\geqslant}\,0$. Then $f$ is either divisible by $x_{i}$ or a polynomial of $x_{j}$ for $j\,{\neq}\,i$. Here the first case does not occur, since $Z$ has non-isolated singularities otherwise. So the assertion follows for $d\,{\geqslant}\,4$.
It remains to show the case d= 3d\,{=}\,3. (Note that the case d= 2d\,{=}\,2 follows from the classification theory of quadratic forms over $\mathbb{C}$}, since these are classified only by the rank.) We assume $Z$ is not a cone. Then all the singularities of $Z$ must be ordinary double points, since d= 3d\,{=}\,3. We may assume $c_{1}\,{=}\,\max_{i}|c_{i}|\,{=}\,1$ changing the order of variables and multiplying $\ell$ by a rational number. Then $f$ cannot contain a monomial of the form $x_{1}^{2}x_{i}$ for any $i\,{\in}\,[1,n]$ (that is, the coefficient of this monomial in $f$ vanishes), since $2\thinspace{+}\thinspace c_{i}\,{=}\,0$. Hence $f^{(1)}$ has an singularity at the origin substituting $x_{1}\,{=}\,1$. Moreover we have
$c_{i}\,{\leqslant}\,0\quad\hbox{for any}\,\,\,i\,{\in}\,[2,n].$
Indeed, $f$ must contain a monomial of the form $x_{1}x_{i}x_{k}$ for some $k\,{\in}\,[1,n]$ in order that the singularity of $f^{(1)}$ is an ordinary double point after substituting $x_{1}\,{=}\,1$. (Indeed, monomials of degree 3 can be neglected for the judgement of whether $\mu\,{=}\,1$ or not, since it is decided by the determinant of the matrix whose $(i,j)$-component is the value at 0 of $\partial_{x_{i}}\partial_{x_{j}}f^{(1)}$ for $i,j\,{\in}\,[2,n]$.) This is however impossible if $c_{i}\,{>}\,0$ (where $1\thinspace{+}\thinspace c_{i}\thinspace{+}\thinspace c_{k}\,{=}\,0$ and $|c_{k}|\,{\leqslant}\,1$).
Since $f$ is irreducible, $f$ must contain a monomial of the form $x_{j}x_{j^{\prime}}x_{j^{\prime\prime}}$ with $j,j^{\prime},j^{\prime\prime}\,{\in}\,[2,n]$ (otherwise $f$ is divisible by $x_{1}$). Here $c_{j},c_{j^{\prime}},c_{j^{\prime\prime}}$ must be 0, since they are nonpositive with $c_{j}\thinspace{+}\thinspace c_{j^{\prime}}\thinspace{+}\thinspace c_{j^{\prime\prime}}\,{=}\,0$. In order that $f^{(1)}$ has an ordinary double point at 0, the polynomial $f$ must contain a monomial of the form $x_{1}x_{j}x_{k}$ for some $k\,{\in}\,[1,n]$ as above. Here $c_{k}$ must be 1-1, since $c_{j}\,{=}\,0$ this time. This $k$ is unique and $c_{i}\,{=}\,0$ for any $i\,{\neq}\,1,k$, applying the above argument to -$\ell$. We may assume k= 2k\,{=}\,2 changing the order of variables. We then see that $f\,{=}\,x_{1}x_{2}\thinspace g\thinspace{+}\thinspace h$ with $g,h$ homogeneous polynomials of $x_{i}$ for $i\,{\in}\,[3,n]$ with degree 1 and 3 respectively. However this cannot give an ordinary double point after substituting $x_{1}\,{=}\,1$ or $x_{2}\,{=}\,1$ (since $n\,{\geqslant}\,4$), where the homogeneous polynomial $h$ of degree 3 can be neglected for the judgement of whether $\mu\,{=}\,1$ or not. So the assertion follows also for d= 3d\,{=}\,3. This finishes the proof of Proposition 3.3. ∎
Remark 3.4.
Let $b_{f}(s),b_{f^{m}}(s)$ be the Bernstein-Sato polynomials of $f$ and $f^{m}$ with $m\,{\in}\,{\mathbb{Z}}_{>0}$. These are polynomials of lowest degrees (or generators of ideals) satisfying the equations
$b_{f}(s)f^{s}=P(s)f^{s+1},\quad b_{f^{m}}(s)(f^{m})^{s}=P^{(m)}(s)(f^{m})^{s+1},$
with $P(s),P^{(m)}(s)\,{\in}\,{\mathcal{D}}_{{\mathbb{C}}^{n},0}[s]$. The first equality implies for $k\,{\in}\,{\mathbb{N}}$ that
$b_{f}(ms\thinspace{+}\thinspace k)f^{ms+k}=P(ms\thinspace{+}\thinspace k)f^{ms+k+1}.$
We then get the divisibility
$\hbox{$\prod$}_{k=0}^{m-1}b_{f}(ms\thinspace{+}\thinspace k)\in b_{f^{m}}(s)\thinspace{\mathbb{Q}}[s].$
On the other hand we see that
$b_{f^{m}}\bigl(\tfrac{s}{m}\bigr)\in b_{f}(s)\thinspace{\mathbb{Q}}[s].$
These show for instance that the maximal root of $b_{f^{m}}(s)$ coincides with that of $b_{f}(s)$ divided by $m$. (This would be quite well known to specialists, see also [Bu 15, Example 4.11].)
Remark 3.5.
For $j\,{\in}\,[1,n]$, let $\eta_{j}$ be the exterior product of the ${\rm d}x_{i}$ for $i\,{\neq}\,j$ up to sign so that ${\rm d}x_{i}{\wedge}\thinspace\eta_{j}\,{=}\,\delta_{i,j}\thinspace\omega$ with $\omega\,{:=}\,{\rm d}x_{1}{\wedge}\cdots{\wedge}\thinspace{\rm d}x_{n}$. The exterior product of ${\rm d}f$ with an (n1)(n{-}1)-form $\hbox{$\sum$}_{i,j}\,c_{i,j}x_{i}\eta_{j}$ for $c_{i,j}\,{\in}\,{\mathbb{C}}$ (which induces the $E_{0}$-differential of the pole order spectral sequence) is equal to $\xi(f)\omega$ with $\xi\,{:=}\,\hbox{$\sum$}_{i,j}\,c_{i,j}x_{i}\partial_{x_{j}}$. The image of the (n1)(n{-}1)-form by the exterior derivative $\rm d$} (which induces the $E_{1}$-differential of the spectral sequence) is given by $\omega$ multiplied by the trace $\hbox{$\sum$}_{i}\,c_{i,i}$, since $\partial_{x_{j}}(c_{i,j}x_{i})\,{=}\,\delta_{i,j}c_{i,j}$. (The first relation is quite implicit in the calculations in [Sa 25, Section 8]. The second seems to have been forgotten to consider.)
Lastly we give an entirely elementary proof of Lemma 1 generalizing an argument in the proof of [Sa 25, Proposition 8.1], see also [BaVe 26, Lemma 3.11].
Proof of Lemma 1.
Let $\xi\,{=}\,\hbox{$\sum$}_{i,j}\,c_{i,j}x_{i}\partial_{x_{j}}$ be a vector field of degree 0 which annihilates a homogeneous polynomial $f\,{=}\,\hbox{$\sum$}_{\nu\in{\mathbb{N}}^{n}}\,a_{\nu}{\bf x}^{\nu}$ of degree $d$. We may assume the matrix $A\,{=}\,(c_{i,j})$ is in Jordan normal form, see [Sa 25, Remark 8.4]. Set $\xi^{\prime}\,{:=}\,\hbox{$\sum$}_{i}\,c_{i,i}x_{i}\partial_{x_{i}}$, $\xi^{\prime\prime}\,{:=}\,\xi\thinspace{-}\thinspace\xi^{\prime}$. Note that $\xi^{\prime\prime}$ is a $\mathbb{C}$}-linear combination of $x_{j}\partial_{x_{j+1}}$ for some $j\,{\in}\,[1,n]$ with $\{j,j{+}1\}\,{\subset}\,I(\alpha^{\prime})$ and $\alpha^{\prime}\,{\in}\,{\mathbb{C}}$, since $A$ is in Jordan normal form.
Put $\ell_{\xi^{\prime}}(\nu)\,{:=}\,\hbox{$\sum$}_{i}\,c_{i,i}\nu_{i}$ for $\nu\,{=}\,(\nu_{1},\dots,\nu_{n})\,{\in}\,{\mathbb{N}}^{n}$. Set $I(\alpha)\,{:=}\,\{i\,|\,c_{i,i}\,{=}\,\alpha\}$ for $\alpha\,{\in}\,{\mathbb{C}}$. Put $\nu(\alpha)_{i}\,{=}\,\nu_{i}$ if $i\,{\in}\,I(\alpha)$ and 0 otherwise. Then ${\bf x}^{\nu}\,{=}\,\hbox{$\prod$}_{\alpha\in{\mathbb{C}}}\,{\bf x}^{\nu(\alpha)}$, that is, $\nu\,{=}\,\hbox{$\sum$}_{\alpha\in{\mathbb{C}}}\,\nu(\alpha)$. This is the decomposition by eigenvalues of $A$. We have
$\ell_{\xi^{\prime}}(\nu)\,{=}\,\hbox{$\sum$}_{\alpha\in{\mathbb{C}}}\,\alpha|\nu(\alpha)|\quad\hbox{with}\quad|\nu(\alpha)|\,{=}\,\hbox{$\sum$}_{i}\,\nu(\alpha)_{i}.$
Set $\Xi_{\xi^{\prime}}\,{:=}\,\ell_{\xi^{\prime}}^{-1}(0)\cap{\mathbb{Z}}^{n}$. We have to show that ${\rm Supp}_{\bf x}f\,{\subset}\,\Xi_{\xi^{\prime}}$, that is $a_{\nu}\,{=}\,0$ if $\nu\,{\notin}\,\Xi_{\xi^{\prime}}$. This can be shown by decreasing induction taking the graded pieces of an increasing filtration $G$ on the vector space ${\mathbb{C}}[{\bf x}]^{d}$ (consisting of homogeneous polynomials of degree $d$) such that the action of the nilpotent part $\xi^{\prime\prime}$ can be neglected. This filtration can be defined so that $G_{p}$ is generated over $\mathbb{C}$} by monomials ${\bf x}^{\nu}$ with $\phi(\nu)\,{\leqslant}\,p$ and $|\nu|\,{=}\,d$, where $\phi(\nu)$ is given for instance by $\hbox{$\sum$}_{i=1}^{n}\,(d{+}1)^{i}(\nu_{i}{+}1)$ on $\{\nu\,{\in}\,{\mathbb{N}}^{n}\,|\,|\nu|\,{=}\,d\}$. Note that
$\xi^{\prime}{\bf x}^{\nu}\,{=}\,\ell_{\xi^{\prime}}(\nu){\bf x}^{\nu}\quad\hbox{for}\,\,\,\,\nu\,{\in}\,{\mathbb{N}}^{n},$
and the monomials ${\bf x}^{\nu}$ are $\mathbb{C}$}-linearly independent in the polynomial ring ${\mathbb{C}}[{\bf x}]$. (This is the key point in the proof.) Here we can verify that for $j\,{\in}\,[1,n]$ with $\{j,j{+}1\}\,{\subset}\,I(\alpha^{\prime})$ and $\alpha^{\prime}\,{\in}\,{\mathbb{C}}$, the subset $\Xi_{\xi^{\prime}}\,{\subset}\,{\mathbb{Z}}^{n}$ is stable  by the self-map $\rho^{(j)}$ of ${\mathbb{Z}}^{n}$ defined by
$\rho^{(j)}(\nu)_{i}\,{=}\,\begin{cases}\nu_{i}\thinspace{+}\thinspace 1&\hbox{if}\,\,\,i\,{=}\,j,\\ \nu_{i}\thinspace{-}\thinspace 1&\hbox{if}\,\,\,i\,{=}\,j\thinspace{+}\thinspace 1,\\ \nu_{i}&\hbox{otherwise,}\end{cases}$
corresponding to $x_{j}\partial_{x_{j+1}}$ (where $\phi\bigl(\rho^{(j)}(\nu)\bigr)\,{<}\,\phi(\nu)$ for $\nu\,{\in}\,{\mathbb{N}}^{n}$), since the $|\nu(\alpha)|$ do not change by the $\rho^{(j)}$. This implies that the influence of the nilpotent part $\xi^{\prime\prime}$ is restricted inside $\hbox{$\sum$}_{\nu\in\Xi_{\xi^{\prime}}\cap{\mathbb{N}}^{n}}{\mathbb{C}}{\bf x}^{\nu}$, since $A$ is in Jordan normal form. So Lemma 1 follows. ∎